OpenStudy (anonymous):

Let a be a rational number with 0

OpenStudy (anonymous):

\[\cos(3 \pi a)+2\cos(2 \pi a)=0\] (Anlge measurents are in radiens ) Prove that \[a=2/3\]

hartnn (hartnn):

did u try writing cos(3pi*a) using the formula of cos 3x ? and cos(2pi*a) using cos 2x and getting a cubic equation ?

hartnn (hartnn):

x=pi*a

OpenStudy (anonymous):

\[4x^3-4x^2-3x+2=0\]

OpenStudy (anonymous):

\[(2x-1)(2x^2-x-2)=0\] \[x=1/2\] \[\cos(3\pi a)=1/2\]

OpenStudy (anonymous):

\[3\pi a=\pm \pi/3+2\pi k\]

OpenStudy (anonymous):

i made mistake \[4x^3+4x^3-3x-2=0\] \[(2x+1)(2x^2+x-2)=0\] \[x=-1/2\] \[\cos(\pi a)=\pm (\frac{ 2\pi }{ 3 })+2\pi k\]

OpenStudy (anonymous):

last step\[\cos (\pi a)=-1/2\] \[\pi a=\frac{ 2 \pi }{ 3 }\]

OpenStudy (anonymous):

yipee

OpenStudy (anonymous):

\[a=2/3\]

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

??

OpenStudy (anonymous):

guys can you say something

hartnn (hartnn):

sorry i was afk, its correct

OpenStudy (anonymous):

thanks